OpenStudy (anonymous):

how much would $500 invested at 5% interest compoundes annually be worth after 10 year?

OpenStudy (anonymous):

\[P=500$,Rate=5% annual, time t=10 years \] \[A=P \left( 1+\frac{ r }{100} \right)^{t}\] \[A=500\left( 1+\frac{ 5 }{ 100} \right)^{10}\] \[A=500\left( \frac{ 21 }{ 20 } \right)^{10}\] you can solve it by calculator or by log tables \[\log A=\log 500+10\left( \log21-\log20 \right)\] then solve taking log and in the end to get A take antilog